# Checklist Engineering Empowers Multilingual LLM Judges

### v1

@@ 1 Introduction
Evaluation is one of the most important tasks in Natural Language Processing (NLP) for assessing a model’s performance on a specific task. Automating this process is highly beneficial and has been considered since the early stages. Traditional evaluation metrics (Papineni et al., 2002) have some drawbacks, such as the necessity of reference answers and a lack of interpretability, which has led to a paradigm shift toward developing Large Language Model (LLM) evaluators, referred to as LLM-as-a-Judge (Gu et al., 2024; Li et al., 2024). These models are also capable of evaluating long-form LLM generations in either a pointwise or pairwise format—meaning grading a single response or selecting the better response out of two, respectively. Some of the advantages of this approach include high adaptability (Bavaresco et al., ), low inference time, and the non-necessity for the evaluated LLM to be active during evaluation (i.e., it does not need to generate further responses). Despite significant efforts to make LLMs multilingual, extending LLM-as-a-Judge to multilingual configurations has received relatively little attention. Although current multilingual LLM judges (Pombal et al., 2025; Doddapaneni et al., 2024b) perform well, their main limitation is their reliance on proprietary models or the need for a large amount of real or synthetic data to fine-tune a capable evaluator, raising concerns regarding cost, time, and efficiency.
Meanwhile, checklists as interpretable evaluation tools (Doddapaneni et al., 2024a; Cook et al., ) are gaining traction for their transparency and structure, although their application to multilingual evaluation remains relatively underexplored. Most existing approaches also lack robust support for pairwise evaluation. For instance, (Wei et al., ) suggests to handle the pairwise setting by selecting the higher-graded response based on independent pointwise scores, which fails to capture the nuanced comparative superiority between responses. In this work, we present CE-Judge, an LLM-as-a-Judge framework that builds and uses engineered checklists for evaluation. It supports multilingual evaluation and both pointwise and pairwise modes. Our pipeline follows a three-stage process: the first two stages generate broad and dynamic checklist items, and the third applies them for judgment. Notably, even when using a lightweight, open-source model without any fine-tuning, our method demonstrates strong performance across different evaluation scenarios.
@@ 2.1 LLM-as-a-Judge
We begin with Zheng et al. (2023), which uses the generative capabilities of LLMs to act as evaluators. These can be grouped into two types: prompt-based and fine-tuned evaluators. For the former, Li et al. (2025) adopts a ”decompose and aggregate” strategy, identifying weighted evaluation aspects and combining their scores to assess candidate responses. For the latter, which has recently gained momentum, Kim et al. (2024b) is a representative work that trains evaluators using large-scale synthetic data in both pointwise and pairwise setups, incorporating a weight-merging technique. In multilingual settings, consistency remains limited, as noted by Fu & Liu (2025). Among the few multilingual methods, Doddapaneni et al. (2024b) translates between languages to anchor outputs in English for consistent scoring, while Pombal et al. (2025) follows Kim et al. (2024b) to generate multilingual evaluation data and fine-tune models accordingly.
@@ 2.2 Checklist-based Evaluators
Several works have explored checklist-based evaluation. RocketEval (Wei et al., ) generates binary checklist items, then optimizes or reweights them to produce final scores. TICK (Cook et al., ) uses instructions to generate checklists, which LLMs use for self-improvement. CheckEval (Lee et al., 2024) defines high-level criteria, then decomposes, diversifies, and filters them to form evaluation checklists. FBI (Doddapaneni et al., 2024a) employs checklists for meta-evaluation to assess evaluator LLMs. Unlike these works, our framework introduces a novel architecture and, notably: (1) extends to multilingual settings; (2) supports pairwise evaluation beyond pointwise framing; and (3) uniquely incorporates broadness, descriptiveness, dynamism, and answer-mentioning in a unified manner.
@@ 3 CE-Judge pipeline
| CE-Judge | 0.77 | 0.81 | 0.77 | 0.72 | 0.78 | 0.77 | 0.75 | 0.84 | 0.78 | 0.76 | 0.78 | 0.77 |
We present our training-free, efficient evaluation framework (Figure 1), consisting of three steps. The pipeline builds an engineered checklist through level-by-level multilingual understanding, integrates input-output linkage for dynamism, and enhances LLM evaluator decisions. All LLM generations are forced in English to leverage its strong performance in that language.
@@ 3.3 Judgment
The final step is judgment. First, the two checklists from the previous steps are concatenated into a unified checklist. It’s important to note that the entire process described so far is for pointwise evaluation. For pairwise evaluation, the process remains the same, except that the previous two steps are applied to two candidate responses instead of one. As a result, we obtain two checklists, one for each candidate. In this step, we provide the untranslated versions of the instruction and response(s) (to avoid translation bias), along with the checklist(s) and the prompt template in A.3. The LLM is then asked to answer a subset of key checklist items and generate evaluation feedback. Unlike prior works where checklist items are marked with ticks, crosses, or weighted scores, here the model exercises discretion in its judgments, and the final evaluation is left to the model’s decision.
@@ 4.1 Experimental setup
In this work, we used the open-source Qwen2.5-7B-Instruct model (Yang et al., 2024) as the backbone LLM, accessed freely via the Novita API222https://novita.ai/. The hyperparameters “temperature”, “top_p”, and “seed” were set to 0, 1, and 42, respectively, to ensure reproducibility. For translation, we employed the free Google Translate API available through the deep-translator Python package333https://deep-translator.readthedocs.io/en/latest/README.html.
@@ 4.2 Datasets
Since our method is training-free, all datasets are used solely for testing. We evaluated our framework in both pointwise and pairwise settings. For pointwise evaluation, we used the student-annotated subset of the LitEval (Zhang et al., 2024), which contains source–target literary translations for four language pairs with human ratings from 1 to 7. For pairwise evaluation, we employed the reasoning and chat subsets of the MM-Eval dataset (Son et al., 2024), covering 11 and 7 languages, respectively. Each input consists of a reasoning question or chat history, with the task being to choose the better of two candidate responses.
@@ 4.3 Evaluation Metrics
| CE-Judge | 0.28 | 0.49 | |||
To evaluate our CE-Judge framework in pointwise mode, we used accuracy—defined as the number of correct predictions over the total number of samples. For the pairwise setting, we measured performance using Kendall’s Tau correlation coefficient (Kendall, 1938), which assesses the agreement between our model’s rankings and human judgments.
@@ 4.4 Baselines
We compare our framework with three types of models. The first includes proprietary models like GPT-4o. The second is Qwen2.5-7B-Instruct, a strong multilingual open-source LLM that is instruction-tuned from a pretrained model without further fine-tuning. The third category consists of models explicitly trained as evaluators, such as Prometheus 2 (Kim et al., 2024b), Hercule (Doddapaneni et al., 2024b), and M-Prometheus (Pombal et al., 2025), as discussed in Subsection 2.1.
@@ 5 Results
We evaluate CE-Judge on multilingual reasoning, chat, and literary translation tasks against proprietary and open-source baselines, including the fine-tuned M-Prometheus.
In the reasoning task (Table 3), CE-Judge achieves the highest average (0.77) among open-source models, outperforming several large fine-tuned evaluators like M-Prometheus 14B (0.70). Despite using only 7B parameters and no training, it performs close to GPT-4o (0.79), excelling in both high- and low-resource languages.
In chat evaluation (Table 4), CE-Judge leads again with 0.75 average accuracy, surpassing GPT-4o (0.73) and significantly outperforming M-Prometheus 7B (0.62).
For literary translation (Table 4.3), which requires fine-grained linguistic and stylistic insight, CE-Judge attains an average Kendall correlation (0.38) comparable to GPT-4o and much better than its backbone model (0.17). It slightly trails M-Prometheus 7B (0.43), which benefited from fine-tuning on large supervised multilingual data.
@@ 6 Conclusion
In this work, we introduce CE-Judge, a novel and straightforward checklist-based framework for multilingual LLM-as-a-Judge that is training-free and built on an open-source model. By leveraging dynamic, broad, and flexible checklist items, CE-Judge supports both pointwise and pairwise evaluations across diverse languages. Experiments on multiple multilingual benchmarks show that CE-Judge not only generally outperforms open-source baselines but also performs comparably to GPT-4o. These results highlight the promise of structured, dynamic evaluation techniques for improving the reliability and interpretability of LLM judgment, particularly in multilingual contexts, for more consistent performance.
7 Future directions
Despite its strong results and training-free design, our framework has several limitations to address in future work. First, crafting effective prompts for each of the three steps per task can be time-consuming, so developing an automatic, adaptable prompt generation module would be beneficial. Second, our method relies solely on LLM generation, which may suffer from misalignment between training objectives and robust text generation. Incorporating internal LLM representations, as shown by Sheng et al. (2024), could capture more accurate implicit knowledge. Finally, our framework’s flexibility suggests potential extensions as a plug-and-play method or adaptations to other evaluation strategies, such as interview-based evaluation (Kim et al., 2024a).

### v2

@@ Abstract
Checklist Engineering Empowers Multilingual LLM Judges
Mohammad Ghiasvand Mohammadkhani Amirkabir University of Technology mohammad.ghiasvand@aut.ac.ir Hamid Beigy Sharif University of Technology beigy@sharif.edu
@@ 1 Introduction
Evaluation is a fundamental task in Natural Language Processing (NLP) for measuring a model’s performance on specific tasks. Automating this process offers significant benefits and has been a focus since the early stages of NLP research. Moreover, beyond creating evaluators proficient in English, it is crucial to develop their evaluation capabilities in parallel for other languages. Traditional evaluation metrics (Papineni et al., 2002) have some drawbacks, such as the necessity of reference answers and a lack of interpretability, which has led to a paradigm shift toward developing Large Language Model (LLM) evaluators, referred to as LLM-as-a-Judge (Gu et al., 2025; Li et al., 2024). These models are also capable of evaluating long-form LLM generations in either a pointwise or pairwise format—meaning grading a single response or selecting the better response out of two, respectively. Some advantages of this approach include high adaptability (Bavaresco et al., ) and interpretability, in contrast to traditional metrics, as well as low inference time and the fact that the evaluated LLM does not need to be active during evaluation (i.e., it does not need to generate additional responses), both in contrast to more complex LLM-based evaluation frameworks such as Kim et al. (2025).
Despite significant efforts to make LLMs multilingual (Qin et al., 2024), extending LLM-as-a-Judge to multilingual configurations has received relatively little attention. Although current multilingual LLM judges (Pombal et al., 2025; Doddapaneni et al., 2025) perform well, their main limitation is their reliance on proprietary models or the need for a large amount of real or synthetic data to fine-tune a capable evaluator, raising concerns regarding cost, time, and efficiency.
Meanwhile, checklists as interpretable evaluation tools (Doddapaneni et al., 2024; Cook et al., ) are gaining traction for their transparency and structure, although their application to multilingual evaluation remains relatively underexplored, and most of them also lack robust support for pairwise evaluation. For instance, (Wei et al., 2025) suggests to handle the pairwise setting by selecting the higher-graded response based on independent pointwise scores, which fails to capture the nuanced comparative superiority between responses. In this work, we present CE-Judge, an LLM-as-a-Judge framework that builds and uses engineered checklists for evaluation. It supports multilingual evaluation and both pointwise and pairwise modes. Our pipeline follows a three-stage process: the first two stages aim to generate broad and dynamic checklist items, and the third applies them for judgment. Notably, by using a lightweight, open-source LLM without any fine-tuning, our method demonstrates strong performance across different evaluation scenarios.
@@ 2.1 LLM-as-a-Judge
We begin with Zheng et al. (2023), which uses the generative capabilities of LLMs to act as evaluators. These can be grouped into two types: prompt-based and fine-tuned evaluators. For instance, for the former, Li et al. (2025) adopts a "decompose and aggregate" strategy, identifying weighted evaluation aspects and combining their scores to assess candidate responses. For the latter, which has recently gained momentum, Kim et al. (2024) is a representative work that trains evaluators using large-scale synthetic data in both pointwise and pairwise setups, incorporating a weight-merging technique. In multilingual settings, consistency remains limited, as noted by Fu and Liu (2025). Among the few multilingual methods, Doddapaneni et al. (2025) translates between languages to anchor outputs in English for consistent scoring, while Pombal et al. (2025) follows Kim et al. (2024) to generate multilingual evaluation data and fine-tune models accordingly. Chang et al. (2025) investigates several aspects of multilingual LLM-based evaluators, including reference-free prompting, the effect of language resource availability, and the impact of fine-tuning. Thellmann et al. (2024) creates various multilingual evaluation benchmarks while exploring the impact of translation and evaluating LLMs.
@@ 2.2 Checklist-based Evaluators
Several works have explored checklist-based evaluation. RocketEval (Wei et al., 2025) generates binary checklist items, then reweights them to produce final scores. TICK (Cook et al., ) uses instructions to generate checklists, which LLMs use for self-improvement. CheckEval (Lee et al., 2024) defines high-level criteria, then decomposes, diversifies, and filters them to form evaluation checklists. FBI (Doddapaneni et al., 2024) employs checklists for meta-evaluation to assess evaluator LLMs. Unlike these works, our framework introduces a novel architecture and, notably: (1) extends to multilingual settings; (2) supports pairwise evaluation beyond pointwise framing; and (3) uniquely incorporates broadness, descriptiveness, dynamism, and answer-mentioning in a unified manner.
@@ 3 CE-Judge pipeline
We present our training-free, efficient evaluation framework (Figure 1), which consists of three steps. The pipeline first constructs an engineered checklist through level-by-level multilingual understanding and integration of input-output linkages to enable dynamism, followed by utilizing this checklist to enhance the decisions of the evaluator LLM. All LLM generations are asked to be in English to leverage its strong performance in the language (Mondshine et al., 2025).
@@ 3.3 Judgment
The final step is judgment. First, the two checklists from the previous steps are concatenated into a unified checklist. It’s important to note that the entire process described so far is for pointwise evaluation. For pairwise evaluation, the process remains the same, except that the previous two steps are applied to two candidate responses instead of one. As a result, after concatenation, we obtain two checklists, one for each candidate. In this step, we provide the untranslated versions of the instruction and response(s)—to avoid translation bias—along with the checklist(s) and the prompt template in A.3. The LLM is then asked to answer a subset of key checklist items and generate evaluation feedback. Unlike prior works where checklist items are marked with ticks, crosses, or weighted scores, here the model exercises discretion in its judgments, and the final evaluation is left to the model’s decision.
@@ 4.1 Experimental setup
In this work, we used the Qwen2.5-7B-Instruct model (Yang et al., 2024) as the backbone LLM, accessed freely via the Novita API222https://novita.ai/. The hyperparameters “temperature”, “top_p”, and “seed” were set to 0, 1, and 42, respectively, to ensure reproducibility. For translation, we employed the free Google Translate API available through the deep-translator Python package333https://deep-translator.readthedocs.io/en/latest/README.html.
@@ 4.2 Datasets
Since our method is training-free, all datasets are used solely for testing. We evaluated our framework in both pointwise and pairwise settings. For pointwise evaluation, we used the student-annotated subset of the LitEval (Zhang et al., 2024), which contains source–target literary translations for four language pairs with human ratings from 1 to 7. For pairwise evaluation, we employed the reasoning and chat subsets of the MM-Eval dataset (Son et al., 2024), covering 11 and 7 languages, respectively. Each input consists of a reasoning question or chat history, with the task being to choose the better of two candidate responses. The reason for utilizing the LitEval and MM-Eval datasets is that the former is one of the only multilingual pointwise evaluation datasets, and the latter is more robust than the well-known M-RewardBench multilingual benchmark (Gureja et al., 2025).
@@ 4.3 Evaluation Metrics
To evaluate our CE-Judge framework in pointwise mode, we measured performance using Kendall’s Tau correlation coefficient (Kendall, 1938), which assesses agreement between our model’s rankings and human judgments. For the pairwise setting, we used accuracy—defined as the number of correct predictions over the total number of samples.
@@ 4.4 Baselines
We compare our framework with three types of models. The first includes proprietary models like GPT-4o. The second is Qwen2.5-7B-Instruct, a strong multilingual open-source LLM that is instruction-tuned from a pretrained model without further fine-tuning. The third category consists of models explicitly trained as evaluators, such as Prometheus 2 (Kim et al., 2024), Hercule (Doddapaneni et al., 2025), and M-Prometheus (Pombal et al., 2025), as discussed in Subsection 2.1.
@@ 5 Results
We evaluate CE-Judge on three multilingual evaluation datasets—reasoning, chat, and literary translation—against proprietary and open-source baselines, including the fine-tuned M-Prometheus. In all three tables, languages are shown by their codes, and, more importantly, the results for the other models are taken from Pombal et al. (2025).
In the reasoning evaluation task (Table 3), CE-Judge achieves an average accuracy of 0.77, outperforming all open-source baselines in all languages, including large fine-tuned evaluators such as M-Prometheus 14B. Despite being training-free and based on a 7B-parameter model, it performs competitively with GPT-4o (which has an average accuracy of 0.79) and maintains strong performance across both high- and low-resource languages.
In the chat evaluation (Table 4), CE-Judge achieves an average accuracy of 0.75, surpassing GPT-4o (with the average of 0.73) and significantly outperforming the M-Prometheus models across nearly all languages. This result highlights the robustness of our checklist-driven approach in conversational scenarios that require nuanced, context-aware judgment.
In the literary translation evaluation (Table 4.3), which requires nuanced linguistic and stylistic understanding, CE-Judge achieves an average Kendall’s Tau correlation of 0.38, significantly outperforming its backbone model, Qwen2.5-7B, and delivering performance comparable to GPT-4o. Although it slightly lags behind M-Prometheus 7B (average of 0.43)—which benefits from fine-tuning on supervised machine translation evaluation data—our training-free approach remains highly competitive.
@@ 6 Conclusion
In this work, we introduce CE-Judge, a novel and straightforward checklist-based framework for multilingual LLM-as-a-Judge that is training-free and built on an open-source model. By leveraging dynamic, broad, and flexible checklist items, CE-Judge supports both pointwise and pairwise evaluations across diverse languages. Experiments on multiple multilingual benchmarks show that CE-Judge not only generally outperforms open-source fine-tuned baselines but also performs on par with GPT-4o. These results highlight the promise of structured, dynamic evaluation techniques for improving the reliability and interpretability of LLM judgment, particularly in multilingual contexts, for more consistent performance.
Ethics Statement
This study aims to advance multilingual evaluation using a training-free approach built on an open-source LLM, prioritizing accessibility and transparency. We leveraged publicly available datasets and APIs, with no collection of personal or sensitive data. All experiments are free from human involvement and pose no privacy or safety risks.
Limitations
Despite its strong results and training-free design, our framework has several limitations to address in future work. First, crafting effective prompts for each of the three steps per task can be time-consuming, so developing an automatic, adaptable prompt generation module would be beneficial. Second, our method relies solely on LLM generation, which may suffer from misalignment between training objectives and robust text generation. Incorporating internal LLM representations, as shown by Sheng et al. (2024), could capture more accurate implicit knowledge. Finally, our framework’s flexibility suggests potential extensions as a plug-and-play method or adaptations to other evaluation strategies, such as interview-based evaluation (Kim et al., 2025).